A beam of parallel light of wavelength 6000 Å is incident normally on a 1 rupee  coin of diameter 3 cm. If an observation screen is placed at a distance of 1 m from  the coin, how many Fresnel zones would be obstructed by it (the coin)? For what  separation between the coin and the screen, only 5 zones would be cut off?


1
Expert's answer
2021-06-08T09:34:16-0400

1)

"m=\frac{d^2}{4\lambda a }=\frac{0.03^2}{4\cdot 6\cdot 10^{-7}\cdot1}=375,"

2)

"a'=\frac{d^2}{4\lambda m'}=\frac{0.03^2}{4\cdot 6\cdot 10^{-7}\cdot 5}=75~\text m."